BCF forecasts a 25% market share of total 1kg blocks of chocolate sold in Otago in December 2019 in Springfield. Reliable indCalculate the sales mix variance, the sales quantity variance and the sales volume variance for each product and in total for the month of December 2019.
Answer #1

Answer :

(a) Sales Mix Varaince = Budgeted Margin ( Actual Quantity - Revised actual Quantity)

Rich Dark = $ 2.50 ( 10800 - 9720 ) = $ 2700 F

Mint Chill = $ 4 ( 3600 - 6480 ) = $ 11520 U

Cranberry Vintage = $ 5 ( 7200 - 5400 ) = $ 9000 F

Total Sales Mix variance = $ 2700 F + $ 11520 U + $ 9000 F = $ 180 F

Working note -1

Computation of Revised actual quantity

Budgeted Quantity Actual Quantity

Revised Actual Quantity

( Actual quantity in Budgeted M)ix

Rich Dark 7200 10800 9720
Mint chill 4800 3600 6480
Cranberray Vintage 4000 7200 5400
Total 16000 21600 21600

F = Favorable

U = Unfavorable

(b) Sales Quantity variance = Budgeted Margin ( Revised Actual Quantity - Budgeted Quantity )

Rich Dark = $ 2.50 ( 9720 - 7200 ) = $ 6300 F

Mint Chill = $ 4 ( 6480 - 4800 ) = $ 6720 F

Cranberry Vintage = $ 5 ( 5400 - 4000) = $ 7000 F

Total Sales Quantity variance = $ 6300 F + $ 6720 F + $ 7000 F = 20020 F

(c) Sales volume variance = Budgeted margin ( Actual Quantity - Budgeted quantity )

Rich Dark = $ 2.50 ( 10800 - 7200 ) = $ 9000 F

Mint Chill = $ 4 ( 3600 - 4800 ) = $ 4800 U

Cranberry Vintage = $ 5 ( 7200- 4000) = $ 16000 F

Total Sales Volume  variance = $ 9000 F + $ 4800 U + $ 16000 F = $ 20200 F